# Visual Haystacks: A Vision-Centric Needle-In-A-Haystack Benchmark

### v2

@@ Acknowledgments
  • Appendix A includes additional analyses of the VHs benchmark, discussing the dataset distribution and key characteristics, along with visual examples of single- and multi-needle tasks.
  • Appendix B describes the implementation of the baseline models used in the VHs experiments and presents comprehensive results from the VHs experiments, including detailed performance metrics for single- and multi-needle tasks.
  • Appendix C explores the potential limitations of the VHs benchmark and the MIRAGE model, along with their societal impacts.
Appendix A More Analyses on VHs Benchmark
In subsection 2.1, we previously discussed the construction of the VHs benchmark dataset. Here, we take a closer look at its key aspects to provide a more detailed understanding. Figure A.1 presents the distribution of anchor and target objects within the VHs dataset. Our dataset exhibits a diverse range of objects, and the question/answer pairs are well-balanced. Additionally, we include a Language-Only baseline in Table C.1 to emphasize that the benchmark task cannot be easily solved without leveraging the visual content. These statistics support that our benchmark is both representative and free from bias.
Furthermore, we visualize real single- and multi-needle examples in Figure A.2 and Figure A.3, respectively. These examples highlight the realism and vision-centric nature of our VHs dataset, establishing it as a robust and reliable visual NIAH benchmark.
Figure A.1: We visualize the distribution of the anchor (needle) and target objects in the VHs dataset, highlighting its diversity and comprehensive coverage of common objects from COCO.
Figure A.2: Real examples of the single-needle VHs challenges (10-image sub-task shown). These visual needles are more realistic and challenging compared to the copy-and-paste approaches in (Reid et al., 2024; Wang et al., 2024b). Models must retrieve the correct image containing the visual needle and answer the associated question. Distractor images are intentionally included to increase the task difficulty.
Figure A.3: Real examples of the multi-needle VHs challenges (10-image subtask shown). This task is significantly more difficult than the single-needle task, as it requires models to integrate information across multiple images in addition to handling distractors.
Appendix B Details on the VHs Experiments
B.1 Baseline Implementation
@@
Caption Aggregation: Another interesting baseline is a non-query-aware approach where we first caption each image and then use an LLM to aggregate these captions to answer the question. In practice, we prompt Qwen2-VL with the instruction, "Please provide detailed and concrete captions of the image," to generate high-quality image descriptions. These captions are then passed to Llama-v3.1-8B, prompted in a specific format (showing in Figure B.1) to integrate the information and provide an answer.
LMMs: For proprietary models, we used the APIs for “gpt-4o-2024-05-13” and “gemini-1.5-pro-001.” For open-source models, we employed their official Huggingface versions, with details provided in Table B.1. Each model was prompted with: “You are given a set of images. Please answer the following question in Yes or No: {real question}.”
@@
Figure B.1: The prompt for Llama-v3.1 (8B) used in the caption aggregation baseline.
@@
Qwen2-VL-7B-Instruct (Wang et al., 2024a) https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct
@@
Table B.1: Open-source models and their corresponding HuggingFace links used in the evaluation.
B.2 Comprehensive Results
@@
In addition to the line charts in Figure 2 and Figure 3, we provide raw data and more baseline comparisons in the appendix. The complete results for the single-needle and multi-needle challenges are presented in Table C.1 and Table C.2, respectively. To better illustrate the positional bias pathology highlighted in Figure 4, we include multiple 1-D plots in Figure C.1 for clearer interpretation. For proprietary models, we observe that Gemini v1.5 Pro tends to favor needle images at the beginning of the sequence, whereas GPT-4o shows a preference for images that are not in the middle. In contrast, for open-source models with a small number of images ($N\leq 5$), LMMs perform significantly better when the image is placed near the question (at the end of the image sequence). However, as the number of images increases, models like Qwen2-VL shift their preference towards images at the beginning, while other models still favor images at the end. We believe this intriguing behavior is influenced by the training data, but further investigation is needed to confirm and address this issue in future work.
Appendix C Potential Limitations and Societal Impact
VHs: While VHs is the first benchmark to investigate how LMMs can retrieve and reason across sets of images, it has some notable limitations. The primary limitation of the benchmark is the scope: since it is based on MS-COCO images, it inherits the implicit biases in the dataset including notable gender, race, and location biases (Hendricks et al., 2018; Bhargava & Forsyth, 2019; Hirota et al., 2022; Wang et al., 2022). It is important to work in the future toward developing benchmarks that do not favor models that prefer data having such biases. Beyond such implicit biases, COCO objects are also limited to 80 categories - strong performance on the VHs benchmark does not imply that the model will generalize well to all MIQA problems. Finally, the VHs benchmark is primarily template-based, which means that it does not evaluate the language reasoning capabilities of the LLM. A more complex benchmark would require making more detailed inferences, and require multi-hop reasoning across a wider range of open-domain object sets.
@@ Appendix C Potential Limitations and Societal Impact
Table C.2: As in the single-needle experiments, the performance of LMMs degrades rapidly as the number of input images (N) increases. The low performance of most LMMs at the beginning (N=5) highlights their difficulty in both retrieving the correct images and integrating information across frames. The definition of "E" is the same as in Table C.1.

### v3

@@ Acknowledgments
  • Appendix A provides information on the code release, including links to the codebases and datasets used in the project.
  • Appendix B includes additional analyses of the VHs benchmark, discussing the dataset distribution and key characteristics, along with visual examples of single- and multi-needle tasks.
  • Appendix C describes the implementation of the baseline models used in the VHs experiments and presents comprehensive results from the VHs experiments, including detailed performance metrics for single- and multi-needle tasks.
  • Appendix D explores the potential limitations of the VHs benchmark and the MIRAGE model, along with their societal impacts.
Appendix A Code Release
The project page of this paper is at: https://visual-haystacks.github.io/. Code for MIRAGE is made publicly available under the MIT license at https://github.com/visual-haystacks/mirage, and is derived from the Apache 2.0-licensed LLaVA codebase (Liu et al., 2023b). Code for the VHs benchmark is made publicly available under the MIT license at https://github.com/visual-haystacks/vhs_benchmark and contains elements drawn from the MIT-Licensed POPE benchmark (Yifan et al., 2023).
Appendix B More Analyses on VHs Benchmark
In subsection 2.1, we previously discussed the construction of the VHs benchmark dataset. Here, we take a closer look at its key aspects to provide a more detailed understanding. Figure B.1 presents the distribution of anchor and target objects within the VHs dataset. Our dataset exhibits a diverse range of objects, and the question/answer pairs are well-balanced. Additionally, we include a Language-Only baseline in Table C.3 to emphasize that the benchmark task cannot be easily solved without leveraging the visual content. These statistics support that our benchmark is both representative and free from bias.
Furthermore, we visualize real single- and multi-needle examples in Figure B.2 and Figure B.3, respectively. These examples highlight the realism and vision-centric nature of our VHs dataset, establishing it as a robust and reliable visual NIAH benchmark.
Figure B.1: We visualize the distribution of the anchor (needle) and target objects in the VHs dataset, highlighting its diversity and comprehensive coverage of common objects from COCO.
Figure B.2: Real examples of the single-needle VHs challenges (10-image sub-task shown). These visual needles are more realistic and challenging compared to the copy-and-paste approaches in (Reid et al., 2024; Wang et al., 2024c). Models must retrieve the correct image containing the visual needle and answer the associated question. Distractor images are intentionally included to increase the task difficulty.
Figure B.3: Real examples of the multi-needle VHs challenges (10-image subtask shown). This task is significantly more difficult than the single-needle task, as it requires models to integrate information across multiple images in addition to handling distractors.
Appendix C Details on the VHs Experiments
C.1 Baseline Implementation
@@
Caption Aggregation: Another interesting baseline is a non-query-aware approach where we first caption each image and then use an LLM to aggregate these captions to answer the question. In practice, we prompt Qwen2-VL with the instruction, "Please provide detailed and concrete captions of the image," to generate high-quality image descriptions. These captions are then passed to Llama-v3.1-8B, prompted in a specific format (showing in Figure C.1) to integrate the information and provide an answer.
LMMs: For proprietary models, we used the APIs for “gpt-4o-2024-05-13” and “gemini-1.5-pro-001.” For open-source models, we employed their official Huggingface versions, with details provided in Table C.1. Each model was prompted with: “You are given a set of images. Please answer the following question in Yes or No: {real question}.”
@@
Figure C.1: The prompt for Llama-v3.1 (8B) used in the caption aggregation baseline.
@@
Qwen2-VL-7B-Instruct (Wang et al., 2024b) https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct
@@
Table C.1: Open-source models and their corresponding HuggingFace links used in the evaluation.
C.2 Comprehensive Results
@@
In addition to the line charts in Figure 2 and Figure 3, we provide qualitative results, raw data, and more baseline comparisons in the appendix. Two failure cases of GPT-4o and Gemini v1.5 Pro are shown in Figure C.2. The complete results for the single-needle and multi-needle challenges are presented in Table C.3 and Table C.4, respectively. To better illustrate the positional bias pathology highlighted in Figure 4, we include multiple 1-D plots in Figure C.4 for clearer interpretation. For proprietary models, we observe that Gemini v1.5 Pro tends to favor needle images at the beginning of the sequence, whereas GPT-4o shows a preference for images that are not in the middle. In contrast, for open-source models with a small number of images ($N\leq 5$), LMMs perform significantly better when the image is placed near the question (at the end of the image sequence). However, as the number of images increases, models like Qwen2-VL shift their preference towards images at the beginning, while other models still favor images at the end. We believe this intriguing behavior is influenced by the training data, but further investigation is needed to confirm and address this issue in future work.
Figure C.2: Qualitative examples on our VHs benchmark between GPT-4o, Gemini-v1.5, and MIRAGE. With a retriever, MIRAGE performs significantly better when dealing with small or non-salient objects.
C.3 Image Token Compression Scheme
As described in subsection 4.1, MIRAGE aims to facilitate future MIQA research with a critical focus on reducing the tokens required per image. To explore this, we evaluated various token compression methods using the LLaVA-v1.5-7B model and their single-image instruction tuning dataset—a simpler setup compared to the full MIRAGE framework with LLama-3.1, a retriever, and our custom data in Table 1. The baseline schemes included max-pooling and average-pooling compression, reducing an image to a single token. While previous work has suggested that Q-Former is sub-optimal (Fan et al., 2024), our results in Table C.2 indicate that Q-Former retains reasonable performance while significantly reducing context length.
Method Tokens/Img VQAv2 GQA Vizwiz TextVQA POPE MMB MMB-CN MM-Vet
Original LLaVA 576 78.5 62.0 50.0 58.2 85.9 64.3 58.3 31.1
3x3 Max-Pooling 64 68.7 56.2 41.3 48.5 83.0 59.2 49.3 24.3
Global Avg. Pooling 1 62.5 51.3 37.7 45.5 79.6 55.0 45.5 18.9
Q-Former (Ours) 32 72.8 56.6 48.0 47.1 83.9 61.5 55.0 27.3
Table C.2: Exploration of various token reduction methods. The Q-former demonstrates the most efficiency in reducing token count while maintaining the majority of general QA performance. All experiments were conducted using the official LLaVA-v1.5-7B model with Vicuna-v1.5 as the LLM and LLaVA’s instruction tuning data. Note that the Q-former variant shows lower performance compared to our final model MIRAGE-8.3B, which leverages better training data and a more powerful LLM, Llama-v3.1-8B.
C.4 Further Analyses on Multi-needle Challenges
As shown in Figure 2 and Figure 3, we found that the performance on the multi-needle track sometimes surpasses that of the single-needle track for image sets larger than 20. We would like to clarify that the performance on the multi-needle track is not necessarily lower than that of the single-needle track. As mentioned in subsection 2.1, the multi-needle track poses questions in the format: “Q: For all images with <anchor object>, do ALL/ANY of them contain <target object>? A: Yes/No.” We can then categorize each data point in the multi-needle track into two cases:
  1. 1.
    The “ANY-Yes” / “ALL-No” QA pairs: These cases are comparatively easier in the multi-needle track since retrieving at least one correct image can suffice. Conversely, the single-needle task demands precise retrieval, posing a greater challenge for models prone to false negatives.
  2. 2.
    The “ANY-No” / “ALL-Yes” QA pairs: These are more difficult in the multi-needle track compared to the single-needle one, as the model must retrieve all relevant images and integrate their information, demanding stronger cross-image reasoning.
The empirical results in Figure C.3 support the above explanation, demonstrating that the benchmark effectively reveals nuanced model behaviors rather than exposing a design flaw. These findings align with those in Figure 3 (A), where LMMs consistently struggle with tasks requiring the integration of information across multiple frames.
Figure C.3: We break down multi-needle challenges to two categories and compare that with the single-needle track. Please refer to subsection C.4 for detailed analyses and discussions.
@@ C.4 Further Analyses on Multi-needle Challenges
Table C.4: As in the multi-needle experiments, the performance of LMMs degrades rapidly as the number of input images (N) increases. The low performance of most LMMs at the beginning (N=5) highlights their difficulty in both retrieving the correct images and integrating information across frames. The definition of "E" is the same as in Table C.3.
@@
Appendix D Potential Limitations and Societal Impact
VHs: Just as the NIAH benchmark (easily solvable with regex) serves as a basic unit test for LLMs in long-context NLP tasks, VHs offers a similarly basic unit test for evaluating LMMs in long-context visual understanding, with the main focus on visual retrieval and reasoning across a large collection of images. While its simplicity and unprecedented scale are key strengths, the primary limitation of the benchmark is the scope: since it is based on MS-COCO images, it inherits the implicit biases in the dataset including notable gender, race, and location biases (Hendricks et al., 2018; Bhargava & Forsyth, 2019; Hirota et al., 2022; Wang et al., 2022). It is important to work in the future toward developing benchmarks that do not favor models that prefer data having such biases. Beyond such implicit biases, COCO objects are also limited to 80 categories - strong performance on the VHs benchmark does not imply that the model will generalize well to all MIQA problems. Finally, the VHs benchmark is primarily template-based, which means that it does not evaluate the language reasoning capabilities of the LLM. A more complex benchmark would require making more detailed inferences, and require multi-hop reasoning across a wider range of open-domain object sets.